Barbara is dragging a 24.0 Kg log up a hill. The hill has a 26.0 degree angle with the horizontal. with what force must she pull in order to make the log accelerate at 2.00 m/s towards the top of the hill
1
Expert's answer
2019-04-10T09:44:55-0400

We have:


"ma=F-mg \sin 26"

"F=m(a+g \sin 26)"

"F=24(2+9.81 \sin 26) =151N."